1. A helium-silver (HgAg) laser has an emission wavelength of 224 nm. Calculate the frequency of this laser emission.

2. Which of the following wavelengths of electromagnetic radiation is the highest in energy?
a. 600 nm
b. 2100 nm
c. 60 nm
d. 300 nm
e. 2100 mm

3. The human eye can detect light at a frequency of 6.12×1014 Hz in the visible region of the spectrum. Calculate the energy, in joules, of an individual photon of this light.
1
Expert's answer
2019-11-06T07:47:27-0500

1.

"ν=\frac{c}{λ} = \frac{2.998\times 10^{8} \frac{m}{s}}{224\times10^{-9}m} =1.34 \times 10^{15} Hz"

2. Energy is inversely proportional to the wavelength


"E=\frac{hc}{λ}"

The lower the wavelength the higher the energy. The lowest wavelength is c. 60 nm

3.


"E = hν = 6.6261\times10^{-34} J\times s \times 6.12\times10^{14} s^{-1} = 4.055\times10^{-19} J"